The perimeter of a triangle is 23 feet. Two sides are equal and the length of the third side is 1 ft less than twice the length of an equal side. What is the length of one of the equal sides?
This is what I did:
2x-1=23
-1
2x=22
divide by 2
x=11
2(11)-1=21
This would make the two equal sides 1ft. Is this right?

Answer by stanbon(75887)

The perimeter of a triangle is 23 feet. Two sides are equal and the length of the third side is 1 ft less than twice the length of an equal side. What is the length of one of the equal sides?
----------------------------------
Equation:
x + x + (2x-1) = 23
4x - 1 = 23
4x = 24
x = 6 ft (length of 2 of the sides)
---
2x-1 = 2*6-1 = 11 ft (length of the 3rd side)
==================================================
cheers,
Stan H.